Twice a certain whole number subtracted from 3 times the square of the number leaves 133. Find the number

Answer by Cromlix(4381)

Hi there,
Twice a certain whole number subtracted
from 3 times the square of the number
leaves 133.
3x^2 - 2x = 133
3x^2 - 2x - 133 = 0
Factorise
(3x + 19)(x - 7)= 0
3x + 19 = 0
3x = -19
x = -19/3
x - 7 = 0
x = 7
so x can either equal
-19/3 or 7
Hope this helps:-)